OpenStudy (valiyuh):

find the midpoint between A(-12,4) and B(6,-8)

OpenStudy (anonymous):

(x1+x2)/2 (y1+y2)/2

OpenStudy (anonymous):

-12+6/2=-6/2=-3 4+-8/2=-4/2=-2 midpoint is (-3,-2)

OpenStudy (valiyuh):

Thank you so much ! I dont know why i was having such a difficult time.

OpenStudy (anonymous):

It's OK...at least you're trying!

OpenStudy (valiyuh):

yes that's true thank you again(: